OpenStudy (anonymous):

The 4th grade class is taking a field trip with 38 people. They use a total of 8 vehicles. Each car holds 4 passengers and each minivan holds 6 passengers. How many of each type of vehicle did they have?

OpenStudy (anonymous):

I will give medals to anyone who types!

OpenStudy (anonymous):

Hello.

OpenStudy (anonymous):

I will help.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Show every step please

OpenStudy (anonymous):

Well you know how many people are going and how many cars they are taking

OpenStudy (anonymous):

yes but, you have to give me the answer and show every single step and I will give a medal and be your fan!

OpenStudy (anonymous):

Luca, don't try to bribe us like that. I'll help you, don't fan or give me a medal unless I get you the right answer.

OpenStudy (anonymous):

Ok -_-

OpenStudy (anonymous):

Well a possible answer could be 5 vans and 2 cars

OpenStudy (anonymous):

Explain AJM

OpenStudy (anonymous):

I know it's just how would you do the equation.

OpenStudy (anonymous):

Well what i did was subtract from 38 and use how many you could fit into a car or van

OpenStudy (anonymous):

x + y=8 4x+6y=38 are the two equations.

OpenStudy (anonymous):

38 8 cars = 16 5 vans = 30 So no AJm

OpenStudy (anonymous):

Oh, that helps a lot luca :D

OpenStudy (anonymous):

np :D

OpenStudy (anonymous):

x + y = 8 4x+6y=38 First solve for x

OpenStudy (anonymous):

please show each step in on comment

OpenStudy (anonymous):

one* sorry

OpenStudy (anonymous):

-6x - 6y = -48 (set the top equation to be multiplied by -6 so y cancels out)

OpenStudy (anonymous):

-6x - 6y = -48 4x + 6y = 38

OpenStudy (anonymous):

well we cant just GIVE you the answer, that would get you nowwhere

OpenStudy (anonymous):

Now you will use vertical addition.

OpenStudy (anonymous):

-2x = -10

OpenStudy (anonymous):

Now you isolate x. x = 5

OpenStudy (anonymous):

Now that we have x, we can solve for y.

OpenStudy (anonymous):

4 + y = 8

OpenStudy (anonymous):

That would be plugging x into the equation

OpenStudy (anonymous):

then isolate y (although it's easy to see what it is already)

OpenStudy (anonymous):

y = 8 - 4

OpenStudy (anonymous):

y = 4

OpenStudy (anonymous):

Oops I plugged it in wrong lol

OpenStudy (anonymous):

x = 5 5 + y = 8 So y = 3

OpenStudy (anonymous):

So 3 vans and 5 cars.

OpenStudy (anonymous):

yeag

OpenStudy (anonymous):

x + y=8 4x+6y=38 X represents cars and Y represents minivans. x + y=8 -x -x Y=8-x 4x+6( 8-x ) =38 4x+48-6x = 38 Add -6x to 4x. -2x+48 = 38 Then subtract 48 to both sides. -2x = -10 Divide -2 to both sides and since the other side is a negative the answer would be a positive. -10/-2 = 5 We don’t know “y” yet so you would subtract what you got for “x” by 8. 8-5=3 X= 5 Y=3 They would need 5 cars and 3 minivans.

OpenStudy (anonymous):

@jcpd910

OpenStudy (anonymous):

is that good?

OpenStudy (anonymous):

Yes, correct.